Problems Find the centroid of the following cross-sections and planes. 6.1 Ref. IGIN 3. 6.2. 7
Determine Iz and I, for the cross-sections in Problems 6.6 through 6.8 and Iz for Problems 6.9 through 6.11. 2 2 2 SU 5 *2
Answer #1

ANSWERS 10 3 62 1 ن 3 وا 63 0 دو ے 3 و 8 = (119) = ) (را) - (3 (, 3) - (59) له : ( ما )) : 3 ا : (درد) - (2.5,9) : (دح ,(7-49 6.2 0 G *** ocal 1x appel + 2373-0232 ai-02 +03 X1=2 이 - CMX 6-94 O2 = 1X2 -2 R2 = 270.5 = 215 a3 = 3x3x6=9 az = 4* -5چاہے امامے (0) <eے 2. ا (3) کے روح 5 دووه 49 A = 6x2 = 2 9 = 2+ 6 + 5 =q A = 286 : 19 ده 2+ 6 5 و 24 Aa: 1972 = 4 9 :16. = 6*23 12 : 4 162 - 2x63 - 36 12 DG3 12423 - 8 12 Hence, Ixx = 14 +12 x (5) 2] + [ 36+12X12] + [8 +24 x34] . 576 Iyy M1.9